In 0.181s, through a distance of 8.05 m , a test pilot’s speed decreases from 88.9 m/s to 0m/s . If the pilot’s mass is 70.0 kg,how much work done against his body?
Answer #1

Work done = force * distance

W = F*d

F = ma

V = u + at

0 = 88.9 + a*0.181

a = -491.16 m/s2

F = ma = 70*491.16 = 34381.2 N

W = FS cos a

= 34381.2 * 8.05* cos 180

= -276768.8 J